OpenStudy (anonymous):

how would you solve sin^2 (pi/4) - cos^2 (pi/3)

myininaya (myininaya):

\[\sin(\frac{\pi}{4})=\frac{\sqrt{2}}{2}, \cos(\frac{\pi}{3})=\frac{1}{2}\]

myininaya (myininaya):

use that just plug in

OpenStudy (anonymous):

one ten or two twenty?

OpenStudy (anonymous):

do what?

myininaya (myininaya):

\[(\sin(\frac{\pi}{4}))^2-(\cos(\frac{\pi}{3}))^2\] is what you have just plug in the numbers I gave you above you can find these numbers on your handy dandy trig cheat cheat

OpenStudy (anonymous):

*sheet

myininaya (myininaya):

shhh lol

OpenStudy (anonymous):

:)

OpenStudy (anonymous):

like that one

myininaya (myininaya):

do you understand to do this: \[(\frac{\sqrt{2}}{2})^2-(\frac{1}{2})^2\]

OpenStudy (anonymous):

so i ignore the square sign, solve for what's inside the parentheses then square the answer?

OpenStudy (anonymous):

oh no! square, then subtract. just like it is written

OpenStudy (anonymous):

so that would be 2/4 - 1/4

OpenStudy (anonymous):

yup

myininaya (myininaya):

yes

OpenStudy (anonymous):

on the 4th post above i ment what u put up earlier